# Exact Coset Sampling forQuantum Lattice Algorithms

### v6

@@ 2.1 Front-end properties and additional conditions
so the outcome probabilities are proportional to the squared magnitude of the inner sum (the global phase $\omega_{M}^{\langle\bm{u},\bm{v}^{\ast}\rangle}$ disappears).
Because $M_{2}=M/2$, we have $\omega_{M}^{M_{2}\langle\bm{u},\bm{k}\rangle}=(-1)^{\langle\bm{u},\bm{k}\rangle}$, which depends only on the parity pattern of $\bm{u}_{[2..n]}$ since $k_{1}=0$ for all $\bm{k}\in\mathcal{K}$. On the event $R$ we have $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$, hence
$\langle\bm{u},\Delta\rangle=2D^{2}\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{2D^{2}Q}=0\pmod{M}$,
and therefore $\omega_{M}^{j\langle\bm{u},\Delta\rangle}=1$ for all $j$. Therefore, conditioned on $R$, the outcome probabilities depend on $\bm{u}$ only through the parity pattern of $\bm{u}_{[2..n]}$.
Now fix any $\bm{v}\in H$. The fibre $F_{\bm{v}}:=\{\bm{u}\in\mathbb{Z}_{M}^{n}:\bm{u}\bmod Q=\bm{v}\}$ consists of the vectors $\bm{u}=\bm{v}+Q\bm{t}$ with $\bm{t}\in\mathbb{Z}_{M/Q}^{n}=\mathbb{Z}_{2D^{2}}^{n}$. Since $Q$ is odd, as $\bm{t}$ ranges over $\mathbb{Z}_{2D^{2}}^{n}$ the parity pattern of $(\bm{v}+Q\bm{t})_{[2..n]}$ is exactly uniform over $\{0,1\}^{n-1}$, independently of $\bm{v}$. Hence each parity pattern occurs in $F_{\bm{v}}$ with the same multiplicity for every $\bm{v}\in H$, and since the conditional probabilities depend only on this parity pattern, $\mu_{E}[F_{\bm{v}}\mid R]$ is constant over $\bm{v}\in H$. This proves uniformity of $\bm{u}\bmod Q$ on $H$ conditioned on $R$. Finally, because $b_{1}=-1$ is a unit in $\mathbb{Z}_{Q}$, $|H|=Q^{n-1}$, so $\Pr[\bm{u}\bmod Q=\bm{0}\mid R]=Q^{-(n-1)}$. ∎
Proposition 2.3 (Canonical-gauge compilation given coherent transcript access).
Assume AC1 and suppose the state-preparation procedure is given as an explicit circuit that admits a deferred-measurement implementation producing a coherent transcript register $E$ from which $v_{1}^{\ast}(E)\bmod M_{2}$ (equivalently, the gauge parameter $\lambda(E)$ defined in AC4) can be computed reversibly. Then one can compile the preparation by appending an $E$-controlled basis translation on $X_{1}$ by a multiple of $D^{2}$ (modulo $M$) so that, conditioned on $E$, the resulting offset satisfies the canonical-gauge requirement $v_{1}^{\ast}(E)\equiv\operatorname{ctr}(X_{1}\bmod D^{2})\pmod{M_{2}}$ (equivalently, $\lambda(E)=0$). If the preparation is a black-box state oracle that does not expose such coherent access to $\lambda(E)$, the canonical-gauge condition is an additional oracle promise and is not implied by AC1 alone.
@@
Assume AC1-AC4. Then the procedure in AC4 defines, for each measured $\tau$, a unitary $U_{\mathrm{corr}}^{(\tau)}$ on the first coordinate register such that $|\varphi_{7}\rangle:=U_{\mathrm{corr}}|\widetilde{\varphi}_{7}\rangle$ has the form (1.2). The quadratic chirp $e^{-2\pi ij^{2}/Q}$ in the loop index is removed on every branch.
@@
Under AC1, we have $X_{1}\equiv-2D^{2}j+v_{1}^{\ast}\pmod{M_{2}}$ with $v_{1}^{\ast}$ independent of $j$. Conditioned on the transcript $E$, the measurement of $\tau:=X_{1}\bmod D^{2}$ is deterministic and yields $\tau\equiv v_{1}^{\ast}\bmod D^{2}$ on every branch. By the gauge choice in AC4, with $\bar{v}_{1}=\operatorname{ctr}(\tau)$ and $v_{1,Q}=\bar{v}_{1}\bmod Q$, we have $(X_{1}\bmod Q)-v_{1,Q}\equiv-2D^{2}j\pmod{Q}$ on every branch. The reversible computation of $j_{x}$ in AC4 is well-defined modulo $Q$ by AC2-AC3. For $x=X_{1}$ on a branch with loop label $j$ we have $j_{x}\equiv j\pmod{Q}$. The unitary $U_{\mathrm{corr}}$ multiplies that branch by $\exp(2\pi ij^{2}/Q)$. It cancels the original chirp $e^{-2\pi ij^{2}/Q}$. The map is diagonal in the computational basis of $X_{1}$. Hence it is unitary. ∎
@@
Additional Condition AC5 isolates the only additional front-end property used by Step $9^{\dagger}$: spectral concentration of the one-dimensional Karst-wave envelope at the resonant frequencies $t/Q$ with $t\equiv 0\pmod{Q}$.
@@
and identify $t(\bm{u})$ with its standard representative in $\{0,\dots,Q-1\}$ when used as an argument of $\widehat{\alpha}_{E}(\cdot)$ on $[0,1)$. Taking the envelope width $\sigma_{J}\gtrsim Q\log n$ yields $|\widehat{\alpha}_{E}(t/Q)|\leq\exp(-\Omega(\log^{2}n))\cdot|\widehat{\alpha}_{E}(0)|$ for all integers $t\not\equiv 0\pmod{Q}$ by standard Poisson-summation estimates. Since $\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}\geq|\widehat{\alpha}_{E}(0)|^{2}$, this implies the leakage bound (2.1).
Remark on linear phases. A time shift $j\mapsto j-j_{0}$ multiplies $\widehat{\alpha}_{E}(\theta)$ by a unit-modulus factor and does not affect its magnitude. In contrast, a nontrivial linear phase $\alpha_{E}(j)\mapsto\alpha_{E}(j)e^{2\pi i\nu j}$ shifts the spectrum $\widehat{\alpha}_{E}(\theta)\mapsto\widehat{\alpha}_{E}(\theta+\nu)$ and can move the dominant residue class away from $t=0$ on the grid $\{t/Q\}$. In our setting, such a run-dependent shift is exactly the residual $\omega_{Q}^{-\lambda(E)j}$ phase arising from a gauge mismatch in AC4; the canonical-gauge promise rules it out.
@@
For the complex-Gaussian Karst-wave front end of chen2024quantum (where, conditioned on a transcript $E$, the post-Step 7 amplitudes factor as $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$), instantiated with any re-tuned configuration that satisfies Definition 2.1, Additional Condition AC5 holds (i.e., the leakage bound (2.1) fails only with probability $\operatorname{negl}(n)$ over the front-end transcript).
@@
Under Chen’s Karst-wave analysis, conditioned on a transcript $E$ the loop envelope $\alpha_{E}(j)$ is (up to a global phase, an integer shift in $j$, and negligible truncation) a discrete Gaussian of width $\sigma_{J}$, and the factorization $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$ holds. An integer shift in $j$ does not affect $|\widehat{\alpha}_{E}(\cdot)|$. Any additional run-dependent linear phase on the $1/Q$ grid would shift the dominant residue class away from $t=0$; in our setting this is precisely the gauge-mismatch phenomenon discussed and ruled out in AC4. Thus it suffices to bound the centered discrete-Gaussian prototype $g_{\sigma}(j):=\exp(-\pi j^{2}/\sigma^{2}).$ Its Fourier series satisfies the Poisson-summation identity
@@
Using $\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}\geq|\widehat{\alpha}_{E}(0)|^{2}$ proves (2.1). ∎
@@ 2.2 Access model
We assume (circuit-level) quantum access to an outcome-conditioned state-preparation procedure that, on each invocation, outputs one copy of the post-Step 7 state $|\widetilde{\varphi}_{7}\rangle$ on coordinate registers $X\in(\mathbb{Z}_{M})^{n}$. We additionally assume a canonical-gauge implementation as in Additional Condition AC4: inside the preparation circuit the first coordinate is translated by an $E$-dependent multiple of $D^{2}$ (modulo $M$) so that, conditioned on $E$, writing $\tau:=X_{1}\bmod D^{2}$ and $\bar{v}_{1}:=\operatorname{ctr}(\tau)$, the offset satisfies $v_{1}^{\ast}(E)\equiv\bar{v}_{1}\pmod{M_{2}}$ (equivalently, $\lambda(E)=0$ in Additional Condition AC4). Consequently, measuring $\tau$ allows the post-processing to compute $v_{1,Q}=\bar{v}_{1}\bmod Q$ and the chirp-correction phase for that run.
For each invocation there exists an unobserved classical transcript $E$ such that, conditioned on $E$, the reduced state on $X$ satisfies Additional Condition AC1 with offset $\bm{v}^{\ast}(E)$ and amplitudes $\alpha_{E}(j,\bm{k})$. Across invocations the transcript is re-sampled, so the offset $\bm{v}^{\ast}(E)$ may change from run to run and cannot be forced to repeat.
@@
for some index set $\mathcal{K}\subseteq\{0\}\times\mathbb{Z}^{n-1}$, where $\Delta=2D^{2}\mathbf{b}$ and the offset $\bm{v}^{\ast}(E)\in(\mathbb{Z}_{M})^{n}$ depends on $E$ but not on $j$. Reducing modulo $M_{2}$ eliminates the $M_{2}\bm{k}$ term and yields
@@
This is immediate from Chen’s post-Step 7 computational-basis form (see the excerpted Eq. (14)): the basis values are $2Dj\,\bm{x}+\bm{v}^{\prime}+\tfrac{M}{2}\bm{k}\bmod M$ with $\bm{k}\in 0\!\mid\!\mathbb{Z}^{n-1}$ and $\bm{x}=D\mathbf{b}$. Substituting $\bm{x}=D\mathbf{b}$ gives $2Dj\,\bm{x}=2D^{2}j\,\mathbf{b}=j\Delta$ with $\Delta=2D^{2}\mathbf{b}$, and writing $\bm{v}^{\ast}(E):=\bm{v}^{\prime}\bmod M$ yields the claimed support in $(\mathbb{Z}_{M})^{n}$. Reducing modulo $M_{2}=M/2$ eliminates the $M_{2}\bm{k}$ term, giving $\mathbf{X}(j)\equiv 2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}(E)\pmod{M_{2}}$. ∎

### v7

@@ 2.1 Front-end properties and additional conditions
so the outcome probabilities are proportional to the squared magnitude of the inner sum.
Because $M_{2}=M/2$, we have $\omega_{M}^{M_{2}\langle\bm{u},\bm{k}\rangle}=(-1)^{\langle\bm{u},\bm{k}\rangle}$, which depends only on the parity pattern of $\bm{u}_{[2..n]}$ since $k_{1}=0$ for all $\bm{k}\in\mathcal{K}$. Let $\bm{u}_{Q}:=\bm{u}\bmod Q\in\mathbb{Z}_{Q}^{n}$ and set $t(\bm{u}):=\langle\mathbf{b},\bm{u}_{Q}\rangle\in\mathbb{Z}_{Q}$. We can reduce $\bm{u}$ modulo $Q$ inside the phase because of the special form of $\Delta$. Write $\bm{u}=\bm{u}_{Q}+Q\bm{t}$ with $\bm{t}\in\mathbb{Z}^{n}$ coordinatewise. Since $\Delta=2D^{2}\mathbf{b}$ and $M=2D^{2}Q$, we have
$\langle Q\bm{t},\Delta\rangle=Q\cdot 2D^{2}\langle\bm{t},\mathbf{b}\rangle=M\cdot\langle\bm{t},\mathbf{b}\rangle\in M\mathbb{Z},$
so $\omega_{M}^{\,j\langle Q\bm{t},\Delta\rangle}=1$ for all integers $j$. Therefore
$\omega_{M}^{\,j\langle\bm{u},\Delta\rangle}=\omega_{M}^{\,j\langle\bm{u}_{Q},\Delta\rangle}=\exp\!\left(2\pi i\,j\,\frac{t(\bm{u})}{Q}\right).$
On the event $R$ we have $t(\bm{u})=0$, hence $\omega_{M}^{j\langle\bm{u},\Delta\rangle}=1$ for all $j$. Therefore, conditioned on $R$, the outcome probabilities depend on $\bm{u}$ only through the parity pattern of $\bm{u}_{[2..n]}$.
Now fix any $\bm{v}\in H$. The fibre $F_{\bm{v}}:=\{\bm{u}\in\mathbb{Z}_{M}^{n}:\bm{u}\bmod Q=\bm{v}\}$ consists of the vectors $\bm{u}=\bm{v}+Q\bm{t}$ with $\bm{t}\in\mathbb{Z}_{M/Q}^{n}=\mathbb{Z}_{2D^{2}}^{n}$. Since $Q$ is odd, as $\bm{t}$ ranges over $\mathbb{Z}_{2D^{2}}^{n}$ the parity pattern of $(\bm{v}+Q\bm{t})_{[2..n]}$ is exactly uniform over $\{0,1\}^{n-1}$, independently of $\bm{v}$. Hence, each parity pattern occurs in $F_{\bm{v}}$ with the same multiplicity for every $\bm{v}\in H$, and since the conditional probabilities depend only on this parity pattern, $\mu_{E}[F_{\bm{v}}\mid R]$ is constant over $\bm{v}\in H$. This proves uniformity of $\bm{u}\bmod Q$ on $H$ conditioned on $R$. Finally, because $b_{1}=-1$ is a unit in $\mathbb{Z}_{Q}$, $|H|=Q^{n-1}$, so $\Pr[\bm{u}\bmod Q=\bm{0}\mid R]=Q^{-(n-1)}$. ∎
Proposition 2.3 (Gauge side information and canonical-gauge compilation).
Assume AC1 and suppose the state-preparation procedure is given as an explicit circuit (or oracle) that outputs the coordinate registers together with a classical side-information register $E$ from which the gauge parameter $\lambda(E)\in\mathbb{Z}_{Q}$ of AC4 (equivalently, $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$) can be obtained. Then Additional Condition AC4 holds: the post-processing can compute $v_{1,Q}$ for the current run and implement the chirp-cancellation unitary on $X_{1}$. Moreover, if $E$ is available coherently during the preparation, one can compile the preparation by appending an $E$-controlled basis translation on $X_{1}$ by $\lambda(E)D^{2}$ (modulo $M$) so that the resulting state satisfies the canonical gauge $\lambda(E)=0$. If the preparation is a black-box state oracle that outputs only the coordinate registers and does not expose $\lambda(E)$ (or $v_{1}^{\ast}(E)\bmod Q$) as side information, then AC4 becomes an additional oracle promise and is not implied by AC1 alone.
@@
Assume AC1-AC4. Then the procedure in AC4 defines, for each run-local residue $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$ (obtainable from $\tau$ and/or transcript side information), a unitary $U_{\mathrm{corr}}^{(v_{1,Q})}$ on the first coordinate register such that $|\varphi_{7}\rangle:=U_{\mathrm{corr}}^{(v_{1,Q})}|\widetilde{\varphi}_{7}\rangle$ has the form Equation (1.2). The quadratic chirp $e^{-2\pi ij^{2}/Q}$ in the loop index is removed on every branch.
@@
Under AC1, we have $X_{1}\equiv-2D^{2}j+v_{1}^{\ast}\pmod{M_{2}}$ with $v_{1}^{\ast}$ independent of $j$. Conditioned on the transcript $E$, the measurement of $\tau:=X_{1}\bmod D^{2}$ is deterministic and yields $\tau\equiv v_{1}^{\ast}\bmod D^{2}$ on every branch. By AC4, the post-processing obtains $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$, hence $(X_{1}\bmod Q)-v_{1,Q}\equiv-2D^{2}j\pmod{Q}$ on every branch. The reversible computation of $j(X_{1})$ in AC4 is well-defined modulo $Q$ by AC2-AC3. For $x=X_{1}$ on a branch with loop label $j$ we have $j(X_{1})\equiv j\pmod{Q}$. The unitary $U_{\mathrm{corr}}$ multiplies that branch by $\exp(2\pi ij^{2}/Q)$. It cancels the original chirp $e^{-2\pi ij^{2}/Q}$. The map is diagonal in the computational basis of $X_{1}$. Hence, it is unitary. ∎
@@
Additional Condition AC5 isolates the only additional front-end property used by Step $9^{\dagger}$: spectral concentration of the one-dimensional Karst-wave envelope on the $Q$-point grid, with dominant mass at the single resonant frequency $t=0$.
@@
and identify $t(\bm{u})$ with its standard representative in $\{0,\dots,Q-1\}$ when used as an argument of $\widehat{\alpha}_{E}(\cdot)$ on $[0,1)$. Taking the envelope width $\sigma_{J}\gtrsim Q\log n$ yields $|\widehat{\alpha}_{E}(t/Q)|\leq\exp(-\Omega(\log^{2}n))\cdot|\widehat{\alpha}_{E}(0)|$ for all integers $t\not\equiv 0\pmod{Q}$ by standard Poisson-summation / discrete-Gaussian Fourier-decay estimates (micciancio2007worst; peikert2010efficient; regev2010learning). Since $\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}\geq|\widehat{\alpha}_{E}(0)|^{2}$, this implies the leakage bound Equation (2.1).
A time shift $j\mapsto j-j_{0}$ multiplies $\widehat{\alpha}_{E}(\theta)$ by a unit-modulus factor and does not affect its magnitude. In contrast, a nontrivial linear phase $\alpha_{E}(j)\mapsto\alpha_{E}(j)e^{2\pi i\nu j}$ shifts the spectrum $\widehat{\alpha}_{E}(\theta)\mapsto\widehat{\alpha}_{E}(\theta+\nu)$ and can move the dominant residue class away from $t=0$ on the grid $\{t/Q\}$. In our setting, if the post-processing were to ignore the gauge term $\lambda(E)D^{2}$ and use $v_{1,Q}=\bar{v}_{1}\bmod Q$ in place of the correct value $v_{1,Q}=(\bar{v}_{1}+\lambda(E)D^{2})\bmod Q$, then the chirp-cancellation unitary would leave a residual linear phase $\omega_{Q}^{-\lambda(E)j}$ (up to a global constant) on the loop index $j$. This would shift the dominant residue class away from $t=0$ on the $Q$-point grid. Additional Condition AC4 (or canonical-gauge compilation) rules out this mismatch by requiring the exact run-local residue $v_{1,Q}$.
@@
For the complex-Gaussian Karst-wave front end of chen2024quantum (where, conditioned on a transcript $E$, the post-Step 7 amplitudes factor as $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$), instantiated with any re-tuned configuration that satisfies Definition 2.1, Additional Condition AC5 holds.
@@
Under Chen’s Karst-wave analysis, conditioned on a transcript $E$ the loop envelope $\alpha_{E}(j)$ is a discrete Gaussian of width $\sigma_{J}$, and the factorization $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$ holds. An integer shift in $j$ does not affect $|\widehat{\alpha}_{E}(\cdot)|$. Any additional run-dependent linear phase on the $1/Q$ grid would shift the dominant residue class away from $t=0$; in our setting, this corresponds to using an incorrect $v_{1,Q}$ in the chirp correction, which AC4 rules out by supplying the run-local gauge value. Thus, it suffices to bound the centered discrete-Gaussian prototype $g_{\sigma}(j):=\exp(-\pi j^{2}/\sigma^{2}).$ Its Fourier series satisfies the Poisson-summation identity (see, e.g., (micciancio2007worst))
@@
Using $\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}\geq|\widehat{\alpha}_{E}(0)|^{2}$ proves Equation (2.1). ∎
@@ 2.2 Access model
We assume circuit-level quantum access to an outcome-conditioned state-preparation procedure that, on each invocation, outputs one copy of the post-Step 7 state $|\widetilde{\varphi}_{7}\rangle$ on coordinate registers $X\in(\mathbb{Z}_{M})^{n}$, together with a classical side-information register $E$ that may include Chen’s measured strings and, for AC4, a short run-local gauge value sufficient to obtain the residue $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$ (equivalently, $\lambda(E)\in\mathbb{Z}_{Q}$). We treat $E$ as a classical output register (i.e., returned in the computational basis). Equivalently, the oracle may measure $E$ before returning it. This ensures that reading $E$ and any deterministic function of $E$ is compatible with maintaining coherence over the loop index $j$ and grid label $\bm{k}$ in $|\widetilde{\varphi}_{7}\rangle$. We do not assume this gauge value is derivable from the Step 7 coordinate registers alone; it is treated as explicit side information returned by the oracle, and since it is fixed conditioned on $E$ it can be measured without disturbing the superposition over $j$ and $\bm{k}$. As a special case, if the known preparation circuit is compiled into the canonical gauge $\lambda(E)=0$ (Proposition 2.3), then $v_{1}^{\ast}(E)\bmod Q$ can be recovered from $\tau=X_{1}\bmod D^{2}$ alone and no explicit transcript access is needed for chirp cancellation.
For each invocation there exists a transcript $E$ such that, conditioned on $E$, the reduced state on $X$ satisfies Additional Condition AC1 with offset $\bm{v}^{\ast}(E)$ and amplitudes $\alpha_{E}(j,\bm{k})$. Across invocations, the transcript is re-sampled, so the offset $\bm{v}^{\ast}(E)$ may change from run to run and cannot be forced to repeat.
@@
for some index set $\mathcal{K}\subseteq 0\!\mid\!\{0,1\}^{n-1}$, where $\Delta=2D^{2}\mathbf{b}$ and the offset $\bm{v}^{\ast}(E)\in(\mathbb{Z}_{M})^{n}$ depends on $E$ but not on $j$. Reducing modulo $M_{2}$ eliminates the $M_{2}\bm{k}$ term and yields
@@
This is immediate from Chen’s post-Step 7 computational-basis form: the basis values are $2Dj\,\bm{x}+\bm{v}^{\prime}+\tfrac{M}{2}\bm{k}\bmod M$ with $\bm{k}\in 0\!\mid\!\mathbb{Z}^{n-1}$ and $\bm{x}=D\mathbf{b}$. Since $M_{2}=M/2$, only $\bm{k}\bmod 2$ affects $\tfrac{M}{2}\bm{k}\bmod M$ (and the phase $i^{\|\bm{k}\|^{2}}$), so one may fix $\bm{k}\in 0\!\mid\!\{0,1\}^{n-1}$ as representatives. Substituting $\bm{x}=D\mathbf{b}$ gives $2Dj\,\bm{x}=2D^{2}j\,\mathbf{b}=j\Delta$ with $\Delta=2D^{2}\mathbf{b}$, and writing $\bm{v}^{\ast}(E):=\bm{v}^{\prime}\bmod M$ yields the claimed support in $(\mathbb{Z}_{M})^{n}$. Reducing modulo $M_{2}=M/2$ eliminates the $M_{2}\bm{k}$ term, giving $\mathbf{X}(j)\equiv 2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}(E)\pmod{M_{2}}$. ∎
2:Oracle: One copy of $|\widetilde{\varphi}_{7}\rangle$ on coordinate registers $X\in(\mathbb{Z}_{M})^{n}$ and side information sufficient to obtain $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$ (Definition 2.7).
3:Output: A reduced sample $\bm{u}_{Q}\in\mathbb{Z}_{Q}^{n}$ (possibly $\bm{0}$).
4:Query the state-preparation oracle to obtain $|\widetilde{\varphi}_{7}\rangle$ on $X$ and run-local side information $E$ (including the gauge value needed for AC4).
5:Measure $\tau\leftarrow X_{1}\bmod D^{2}$ and set $\bar{v}_{1}\leftarrow\operatorname{ctr}(\tau)$ $\triangleright$ Deterministic conditioned on $E$ (AC1); needed only if $E$ encodes $\lambda(E)$.
6:Using $(E,\bar{v}_{1})$, compute the run-local residue $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$ as in AC4:
7:   either read $v_{1,Q}$ directly from $E$, or read $\lambda(E)\in\mathbb{Z}_{Q}$ from $E$ and set $v_{1,Q}\leftarrow(\bar{v}_{1}+\lambda(E)D^{2})\bmod Q$.
8:Apply the diagonal unitary $U_{\mathrm{corr}}^{(v_{1,Q})}$ on $X_{1}$ as defined in AC4 (using $\mathrm{inv}$).
9:Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to $X$ and measure $\bm{u}\in\mathbb{Z}_{M}^{n}$.
10:return $\bm{u}_{Q}\leftarrow\bm{u}\bmod Q$.